# Local Dissipativity Analysis of Nonlinear Systems

### v1

@@ 6 Numerical examples
We demonstrate the efficacy of our method on three systems. We first find local conic bounds for a one-dimensional nonlinear system and compare to bounds found through ad-hoc ingenuity. Next, we determine the local $\mathcal{L}_{2}$ gain of a two-dimensional pendulum. Finally, we compare our results to that of [summers2013quantitative] for a three-dimensional polynomial system.
@@ 6.2 $\mathcal{L}_{2}$ gain of Pendulum
where $x\in\mathbb{R}^{n}$, $u\in\mathbb{R}$, and $y\in\mathbb{R}$. Let $\mathcal{X}\supseteq\Omega$ be the triangulated region about the origin, shown in Figure 3, where the RPI set $\Omega$ is found using [lavaei2025L2, Theorem 5], where the system is able to have inputs of $\left\lvert u\right\rvert<0.1942$. Problem 4.2 was applied to system – solving for the $\mathcal{L}_{2}$ gain specifically. Because the pendulum has a sinusoidal term, SOS techniques [summers2013quantitative] cannot be used for analysis. Instead, we compare to the gain found using the global analytical Lyapunov function $V(x)=(1-\cos x_{1})+\frac{1}{2}x_{2}^{2}$, which results in the analytical gain bound $\gamma\geq 1.$ Figure 2 shows the gain found on $\Omega$ for an increasing number of simplices, where the best result is $\gamma=2.221$ Figure 4 shows an example of a storage function found across $\Omega$ with 652 simplices.
@@ 7 Conclusion
This paper presents a general method to determine the local dissipativity properties of nonlinear control affine systems. We develop a convex optimization program that simultaneously optimizes the desired IO characterization of the system and synthesizes the corresponding storage function. Further, we provide guarantees that our analysis method will always be feasible when determining CPA or quadratic storage functions for systems with strict local dissipativity properties.
@@ Appendix A Function Error Bounds
Consider $\zeta:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m},$ where $\zeta\in\mathcal{C}^{2}$. Let $\sigma\vcentcolon=\text{co}\{\mathbf{x}_{j}\}_{j=0}^{n}$ be an n-simplex in $\mathbb{R}^{n}.$ If $\sum_{j=0}^{n}\lambda_{j}x_{j}\in\sigma$, then
@@
By definition, any point $x\in\sigma$ can be written as a convex combination of the vertices, i.e., $x=\sum_{j=0}^{n}\lambda_{j}x_{j}.$ Let $\tilde{e}(\mathbf{x})=\zeta(x)-\sum_{j=0}^{n}\lambda_{j}\zeta(x_{j}).$ Consider a dimension of $\zeta$, represented as $\zeta^{(k)}$, and apply Theorem 8 about $x$ for each vertex point in $\tilde{e}(\mathbf{x})$ to get
$\tilde{e}(x)=\zeta(x)-\sum_{j=0}^{n}\lambda_{j}$ $\Big(\zeta^{(k)}(x)+\langle\nabla\zeta^{(k)}(x),\Delta\mathbf{x}_{j}\rangle$
@@
In each dimension, the zero-th order term of each function’s Taylor expansion cancel with $\zeta^{(k)}(\mathbf{x})$ and the summed first order terms are zero as $\sum_{j=0}^{n}\lambda_{j}\Delta\mathbf{x}_{j}=\sum_{j=0}^{n}\lambda_{j}\mathbf{x}_{j}-\mathbf{x}=0.$ Altogether,
$\tilde{e}(\mathbf{x})=-\frac{1}{2}\sum_{j=0}^{n}\lambda_{j}\begin{bmatrix}\overline{\mathbf{H}}_{\zeta^{(1)}}(x_{j},x)\\ \vdots\\ \overline{\mathbf{H}}_{\zeta^{(m)}}(x_{j},x)\end{bmatrix}=-\frac{1}{2}\sum_{j=0}^{n}\lambda_{j}\overline{\mathbf{H}}_{\zeta}(x_{j},x).$
To find (31), recall that $\left\lVert\tilde{e}(\mathbf{x})\right\rVert_{2}\leq\left\lVert\tilde{e}(\mathbf{x})\right\rVert_{1}.$ Therefore,
$\left\lVert\tilde{e}(\mathbf{x})\right\rVert_{2}\leq\frac{1}{2}\sum_{j=0}^{n}\lambda_{j}\sum_{k=1}^{m}\left\lvert\overline{\mathbf{H}}_{\zeta^{(k)}}(x_{j},x)\right\rvert$.
@@
Then, [giesl2012construction, Lemma 2.3], and the bound $\left\lVert x_{j}-x\right\rVert_{2}^{2}\leq\max_{\nu\in\mathbb{Z}_{0}^{n}}\left\lVert\mathbf{x}_{j}-\mathbf{x}_{\nu}\right\rVert_{2}^{2}$ are applied to find (31).
@@
$\left\lVert\tilde{e}(\mathbf{x})\right\rVert_{\infty}=\frac{1}{2}\sum_{j=0}^{n}\lambda_{j}\max_{k\in\mathbb{Z}_{1}^{m}}\left\lvert\overline{\mathbf{H}}_{\zeta^{(k)}}(x_{j},x)\right\rvert$
@@ Appendix C Analytical Conic Bound
$(Cx+Du)^{\top}\mathbf{Q}(Cx+Du)+2(Cx+Du)^{\top}\mathbf{S}u+u^{\top}\mathbf{R}u,$

### v2

@@ 6 Numerical examples
We demonstrate the efficacy of our method on three systems. We first find local conic bounds for a one-dimensional nonlinear system. Here, we must compare to bounds found through ad-hoc ingenuity because we know of no existing, similar technique to systematically seek generic conic bounds of nonlinear systems. Next, we determine the local $\mathcal{L}_{2}$ gain of a two-dimensional pendulum. Finally, we compare our results to that of [summers2013quantitative] for a three-dimensional polynomial system.
@@ 6.2 $\mathcal{L}_{2}$ gain of Pendulum
where $x\in\mathbb{R}^{n}$, $u\in\mathbb{R}$, and $y\in\mathbb{R}$. Let $\mathcal{X}\supseteq\Omega$ be the triangulated region about the origin, shown in Figure 3, where the RPI set $\Omega$ is found using [lavaei2025L2, Theorem 5], where the system is able to have inputs of $\left\lvert u\right\rvert<0.1942$. Problem 4.2 was applied to the system – solving for the $\mathcal{L}_{2}$ gain specifically. Because the pendulum has a sinusoidal term, SOS techniques [summers2013quantitative] cannot be used for analysis. Instead, we compare to the gain found using the global analytical Lyapunov function $V(x)=(1-\cos x_{1})+\frac{1}{2}x_{2}^{2}$, which results in the analytical gain bound $\gamma\geq 1.$ Figure 2 shows the gain found on $\Omega$ for an increasing number of simplices, where the best result is $\gamma=2.221$ Figure 4 shows an example of a storage function found across $\Omega$ with 652 simplices.
@@ 7 Conclusion
This paper presents a general method to determine the local dissipativity properties of nonlinear control affine systems. We develop a convex optimization program that simultaneously optimizes the desired IO characterization of the system and synthesizes the corresponding storage function. Further, we provide guarantees that our analysis method will always be feasible when determining CPA or quadratic storage functions for systems with strict local dissipativity properties. Prior results establishing dissipativity characterizations are either limited to polynomial systems, only consider gain, or rely on ad hoc, ingenuity. Numerical results show that our techniques produce similarly tight characterizations to alternatives (when those alternatives are available).
@@ Appendix A Function Error Bounds
Consider $\zeta:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m},$ where $\zeta\in\mathcal{C}^{2}$. Let $\sigma\vcentcolon=\text{co}\{x_{j}\}_{j=0}^{n}$ be an n-simplex in $\mathbb{R}^{n}.$ If $\sum_{j=0}^{n}\lambda_{j}x_{j}\in\sigma$, then
@@
By definition, any point $x\in\sigma$ can be written as a convex combination of the vertices, i.e., $x=\sum_{j=0}^{n}\lambda_{j}x_{j}.$ Let $\tilde{e}(x)=\zeta(x)-\sum_{j=0}^{n}\lambda_{j}\zeta(x_{j}).$ Consider a dimension of $\zeta$, represented as $\zeta^{(k)}$, and apply Theorem 8 about $x$ for each vertex point in $\tilde{e}(x)$ to get
$\tilde{e}(x)=\zeta(x)-\sum_{j=0}^{n}\lambda_{j}$ $\Big(\zeta^{(k)}(x)+\langle\nabla\zeta^{(k)}(x),\Delta x_{j}\rangle$
@@
In each dimension, the zero-th order term of each function’s Taylor expansion cancel with $\zeta^{(k)}(x)$ and the summed first order terms are zero as $\sum_{j=0}^{n}\lambda_{j}\Delta x_{j}=\sum_{j=0}^{n}\lambda_{j}x_{j}-x=0.$ Altogether,
$\tilde{e}(x)=-\frac{1}{2}\sum_{j=0}^{n}\lambda_{j}\begin{bmatrix}\overline{\mathbf{H}}_{\zeta^{(1)}}(x_{j},x)\\ \vdots\\ \overline{\mathbf{H}}_{\zeta^{(m)}}(x_{j},x)\end{bmatrix}=-\frac{1}{2}\sum_{j=0}^{n}\lambda_{j}\overline{\mathbf{H}}_{\zeta}(x_{j},x).$
To find (30), recall that $\left\lVert\tilde{e}(x)\right\rVert_{2}\leq\left\lVert\tilde{e}(x)\right\rVert_{1}.$ Therefore,
$\left\lVert\tilde{e}(x)\right\rVert_{2}\leq\frac{1}{2}\sum_{j=0}^{n}\lambda_{j}\sum_{k=1}^{m}\left\lvert\overline{\mathbf{H}}_{\zeta^{(k)}}(x_{j},x)\right\rvert$.
@@
Then, [giesl2012construction, Lemma 2.3], and the bound $\left\lVert x_{j}-x\right\rVert_{2}^{2}\leq\max_{\nu\in\mathbb{Z}_{0}^{n}}\left\lVert x_{j}-x_{\nu}\right\rVert_{2}^{2}$ are applied to find (30).
@@
$\left\lVert\tilde{e}(x)\right\rVert_{\infty}=\frac{1}{2}\sum_{j=0}^{n}\lambda_{j}\max_{k\in\mathbb{Z}_{1}^{m}}\left\lvert\overline{\mathbf{H}}_{\zeta^{(k)}}(x_{j},x)\right\rvert$
@@ Appendix C Analytical Conic Bound
$(Cx+Du)^{\top}\mathbf{Q}(Cx+Du)+2(Cx+Du)^{\top}\mathbf{S}u$
$+u^{\top}\mathbf{R}u,$